At a price of $3.10 a gallon, the weekly demand for gas is 42 gallons. If the price rises to $3.25 a gallon, the demand drops to 39 gallons. Find the formula for quantity of gas demanded in terms of “p”, the price per gallon, assuming the demand is linear.
Answer by Aldorozos(172)

Equation of a line when we have two points on the line: y-y1 = [(x2-x1)/(y2-y1]/(x-x1)
http://www.coolmath.com/algebra/08-lines/12-finding-equation-two-points-01

y-42 = [(3.25 -3.10)/39-42](x-3.1)
y-42 = .15/(-3) *(x-3.1)
y-43 = -.05(x-3.1)This is an equation of a line.